OpenStudy (anonymous):

Been stuck on this for 2 days now can somone help me!!! Part 1. Using the two functions listed below, insert numbers in place of the letters a, b, c, and d so that f(x) and g(x) are inverses. f(x)= x+a b g(x)=cx−d Part 2. Show your work to prove that the inverse of f(x) is g(x). Part 3. Show your work to evaluate g(f(x)). Part 4. Graph your two functions on a coordinate plane. Include a table of values for each function. Include 5 values for each function. Graph the line y = x on the same graph @ganeshie8

OpenStudy (anonymous):

@iGreen @itsbribro @AriPotta

OpenStudy (anonymous):

btw its x+a/b

OpenStudy (anonymous):

Anyone??

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

i wil try to ask somebody

OpenStudy (itsbribro):

Okay for inverse functions, you switch the places of the x and y and solve for y, or in this case x and f(x) and solve for f(x).

OpenStudy (anonymous):

but i have to plug in numbers it says for all of them. abcd

OpenStudy (alexandervonhumboldt2):

ia sked @perl

OpenStudy (anonymous):

he was in here earlier and he left.

OpenStudy (itsbribro):

okayy

OpenStudy (itsbribro):

"Using the two functions listed below, insert numbers in place of the letters a, b, c, and d so that f(x) and g(x) are inverses." So let's make f(x)=x+3/5 and g(x)=5x-3 They're still inverses, we just put numbers in like they asked. Make sense?

OpenStudy (itsbribro):

does that kinda make sense????

OpenStudy (anonymous):

yes

OpenStudy (itsbribro):

for part 1

OpenStudy (anonymous):

okay so thats part 1?

OpenStudy (itsbribro):

yeah

OpenStudy (perl):

are you sure that f(x)=x+3/5 and g(x)=5x-3 are inverses?

OpenStudy (perl):

Check to see if they are inverses does f(g(x)) = x f(g(x)) = f ( 5x - 3) = (5x-3) + 3/5 , which does not equal to x

OpenStudy (anonymous):

okay so how do you make them equal x?

OpenStudy (itsbribro):

ohhhhh oops

OpenStudy (perl):

Is f(x) = x + a/b , and g(x) = cx -d did i copy this correctly?

OpenStudy (anonymous):

yes

OpenStudy (perl):

the easiest inverses are f(x) = x + 1 , g(x) = x- 1

OpenStudy (anonymous):

but its f(x)=x+1/?

OpenStudy (anonymous):

and there is a cx not just x unless c is 1

OpenStudy (perl):

1 = 1/1

OpenStudy (perl):

To make it more interesting f(x) = x + 2 , g(x) = x -2 and you can use 2 = 6/3 , f(x) = x + 6/3, g(x) = 1*x - 2 , that will work

OpenStudy (anonymous):

so that is part 1 f(x)=x+6/3, and g(x)=1*x-2?

OpenStudy (perl):

yes that works

OpenStudy (perl):

now for part 2) you have to prove they are inverses

OpenStudy (anonymous):

and they have to undo each other right?

OpenStudy (perl):

right

OpenStudy (anonymous):

okay so how do i prove they are inverses

OpenStudy (perl):

the directions say exactly Part 2. Show your work to prove that the inverse of f(x) is g(x).

OpenStudy (perl):

it might help to show with a diagram

OpenStudy (anonymous):

okay so what do i show?

OpenStudy (perl):

my draw button died, one sec

OpenStudy (anonymous):

@perl

OpenStudy (perl):